Dixie Bank offers a certificate of deposit with an option to select your own investment period. Jonathan has $7,000 for his CD investment. If the bank is offering a 6​% interest​ rate, compounded​ annually, how much will the CD be worth at maturity if Jonathan picks a two year investment​ period? a five​-year investment​ period? a eight​-year investment​ period?

a fifteen year year investment​ period?

a. How much will the $7,000 CD investment at 6​% interest rate be worth at maturity if Jonathan picks a 2​-year investment​ period?

​(Round to the nearest​ cent.)
Answer #1

a.

Future Value after 2 year investment period = 7,000(1.06)2

Future Value after 2 year investment period = $7,865.20

b.

Future Value after 5 year investment period = 7,000(1.06)5

Future Value after 5 year investment period = $9,367.58

c.

Future Value after 8 year investment period = 7,000(1.06)8

Future Value after 8 year investment period = $11,156.94

d.

Future Value after 15 year investment period = 7,000(1.06)15

Future Value after 15 year investment period = $16,775.91